OpenStudy (anonymous):

What is the remainder when x^2+4 is divided by x-2 ?

OpenStudy (anonymous):

whatever you get when you replace x by 2 in the original

OpenStudy (anonymous):

i get \[2^2+4=8\]

OpenStudy (anonymous):

I would get 8 also :D

OpenStudy (anonymous):

\[\frac{x^2+4}{x-2}=x+2+\frac{8}{x-2} \]